OpenStudy (coolio100):

At a temperature of t degree Fahrenheit, sound travels S feet per second, where S=21.9sqrt(5t+2457). During blasting for avalanche control in the mountains, sound traveled at a rate of 1110 ft/sec. what is the temperature

OpenStudy (brokenstanzas):

Well, I believe S would equal 1110

OpenStudy (brokenstanzas):

Can you work it out from there?

OpenStudy (coolio100):

@brokenstanzas I'm just confused how To get the t out of the square root cause I'm pretty sure I have to solve for that not S.

OpenStudy (coolio100):

This is what it looks like for reference

OpenStudy (brokenstanzas):

That would be the next step after substituting 1110 for S since you can get much of an answer if you try to solve it with two variables. After the substitution, you would square \[\sqrt{5t+2457}\] and solve from there

OpenStudy (coolio100):

@brokenstanzas so I got 1110=21.9(5t)+2457

OpenStudy (coolio100):

What would the answer be I'm still stuck

OpenStudy (sshayer):

\[\frac{ 1110 }{ 21.9 }=\sqrt{5t+2457}\] \[\left( \frac{ 11100 }{ 219 } \right)^2=5t+2457\] \[\left( \frac{ 3700 }{ 73 } \right)^2-2457=5t\] find t

OpenStudy (coolio100):

@sshayer I got 22 as my final answer